3.1 Find the p 50, x-: 9.5, s 7. In a test of Ho: μ vlaue. (ans. 0.1271) 10 vs. Ha: μ < 10, for a sample with n
Answer #1

Solution:

To test the hypothesis,

Ho: \mu = 10 vs H1: \mu\neq 10

n = 50

\bar x = 9.5

s = 3.1

First we find the test statistic.

t =   \frac{\bar x - \mu}{s/\sqrt{n}} = 9.5 10 3.1/50

t = -1.14

Now,since sample size is large(greater than 30),so we can use z table to find P value.

First observe our H1. Since there is < sign, so the test is one sided left tailed.

so,P value = P(Z < -1.14) = 0.1271 (using z table